# Modeling AdaGrad, RMSProp, and Adam with Integro-Differential Equations

### v1

@@ 4 Numerical Simulations
In this section, we present numerical simulations [^3] All simulations can be reproduced from the repository available at GitHub. that validate the approximations proposed in the preceding sections. As shown, the simulations closely mirror the behavior of the discrete algorithms, confirming that the first-order integro-differential equations serve as accurate and reliable models for analyzing the dynamics of these optimization methods. Additional numerical simulations related to the Mean Squared Error (MSE) loss function are provided in Appendix B. These further demonstrate, with an alternative example, how continuous models can effectively capture the dynamics of these optimization algorithms. Specifically, we simulate the convergence of the loss function while approximating the target function $\hat{y}(x)=2x$ using the model function $y(x)=\theta x$.
@@ 4.1 Numerical Method
To solve these equations, we employed the IDESolver method as described in Gelmi & Jorquera (2014), which we implemented from scratch, tailoring the method to fit our specific requirements rather than using the existing Python module. See the pseudocode in Algorithm 4.
@@
For the integral calculations, we employed Gaussian quadrature (fixed_quad) from SciPy with $n=1000$, in place of the quad method used in the original paper. This leads to a significant increase in calculation speed. For the Adam optimizer, we further enhanced computational efficiency by parallelizing the integral computations.
Additionally, we used the Euler method to solve the differential equation, rather than the RK45 method commonly employed in the original approach, to ensure consistency with Assumption (2). The (future) time increment term $\alpha$ of Propositions (1) and (2) is obtained using the interp1d interpolation function from SciPy.
@@
[…]
@@
[…]
@@ 4.2 Numerical Results
In this section, we present the numerical results obtained using the numerical method described previously and compare them with the outcomes derived from the original algorithms.
@@
Figure 8 shows the convergence of $\theta$ values using the nonlocal continuous AdaGrad method for minimizing the function $(x-4)^{2}$, evaluated at two learning rates: $0.1$ and $0.01$. This figure demonstrates that the nonlocal continuous AdaGrad method effectively replicates the optimization dynamics of the conventional AdaGrad method.
Figure 8: Convergence of $\theta(t)$ using the first-order nonlocal continuous AdaGrad method. The plot illustrates the convergence trajectories for minimizing the function $(x-4)^{2}$ using two different learning rates: $0.1$ (left) and $0.01$ (right). At a higher learning rate ($0.1$), the algorithm rapidly converges to the target value $\theta=4$, stabilizing in under 1,500 k-iterations. With a lower learning rate ($0.01$), the convergence is more gradual, reaching the target around 100,000 k-iterations. To clearly demonstrate how accurately the nonlocal continuous model reproduces the discrete dynamics, the plot also includes the simulation of the discrete AdaGrad, making it easier to observe their alignment.
@@
For $\theta$, with a learning rate of 0.1, $\beta=0.99$ leads to faster convergence, while $\beta=0.9$ and $\beta=0.0$ result in slower convergence, with the latter exhibiting more linear behavior but showing noticeable oscillations. When the learning rate is reduced to 0.01, the overall convergence slows, and the oscillatory behavior for $\beta=0.0$ becomes less pronounced [^4] A slight oscillation is observed for $\beta=0.9$ and a learning rate of 0.01, induced by the numerical method, as the convergence error does not decrease below $1\times 10^{-4}$..
Figure 10: Convergence of $\theta(t)$ using the first-order nonlocal continuous RMSProp method. The plot shows the convergence to the minimum value of $\theta=4$ for the function $(\theta-4)^{2}$. The left subplot corresponds to a learning rate of 0.1, while the right subplot uses a learning rate of 0.01. With a learning rate of 0.1, $\beta=0.0$ exhibits more noticeable oscillations, whereas $\beta=0.9$ and $0.99$ converge more smoothly. For a learning rate of 0.01, a slight destabilization occurs for $\beta=0.9$ as the solution approaches the final result, caused by the numerical method. We have also included the discrete RMSProp simulation to facilitate comparison. At higher learning rates, slight differences are observed: for $\beta=0.0$, the oscillations in the discrete case start immediately upon reaching the minimum, whereas in the continuous case, they take a few k-iterations to begin. On the other hand, for $\beta=0.99$, the curve is more pronounced in the discrete model compared to the continuous one.
@@
Figure 11: Convergence trajectories of $v(t)$ for the first-order nonlocal continuous RMSProp optimizer. This plot illustrates the convergence of the squared gradient moving average $v(t)$. The left subplot represents a learning rate of 0.1, while the right subplot corresponds to a learning rate of 0.01. For $\beta$ values of 0.0 and 0.9, a slight initial bump is noticeable, but both eventually decay towards zero. To enable a direct comparison, we have also included the discrete RMSProp case, demonstrating the strong alignment between the nonlocal continuous model and its discrete counterpart. The main differences appear with a learning rate of 0.1, where the values are slightly higher for $\beta=0.9$ and $\beta=0.99$, and for $\beta=0.0$, there is a slight bump rather than a direct descent.
When compared to the standard RMSProp optimizer, the nonlocal continuous variant exhibits a similar overall convergence pattern but with subtle differences. In the $\theta$-plot for a learning rate of 0.1 and $\beta=0.0$, stability is achieved, but oscillations begin after a few iterations, in contrast to the conventional case where oscillations start immediately. Additionally, in the squared gradients plot, for $\beta=0.0$, the continuous case initially shows a slight bump before gradually decaying, whereas in the discrete case, the decay occurs immediately without any “initial” increase. This difference arises because the continuous simulation captures a smoother temporal progression, while the discrete method reflects a more abrupt change.
Furthermore, for $\beta=0.9$ and $\beta=0.99$, the nonlocal continuous model displays a slightly different dynamic compared to the discrete case, reaching slightly higher values. However, as the learning rate decreases, these differences diminish, and the models become almost identical, showcasing a near-perfect accuracy in the continuous representation.
@@
Figures 12, 13, and 14 illustrate the behavior of the first-order nonlocal continuous Adam optimization model by showing the convergence trajectories of $\theta(t)$, $m(t)$, and $v(t)$ over time for different parameter settings ($\beta_{1}$ and $\beta_{2}$) and learning rates ($\alpha=0.1$ and $\alpha=0.01$).
@@
When compared to the discrete Adam model, the nonlocal continuous model is nearly identical at lower learning rates. However, for a learning rate of $0.1$, slight differences can be observed: initially, for $\beta_{1}=0.0$, the continuous model diverges slightly, and for $\beta_{1}=0.9$, the peak is marginally higher than in the simulated discrete case. Nonetheless, these differences are minimal. As the learning rate decreases, both methods become almost exactly the same.
Figure 12: Convergence of $\theta(t)$ using the first-order nonlocal continuous Adam method. The plot illustrates the convergence trajectories of $\theta$-values for the first-order nonlocal continuous Adam model under different parameter settings of $\beta_{1}$ and $\beta_{2}$ with two distinct learning rates ($0.1$ and $0.01$). For $\beta_{1}=0.9$ and a learning rate of 0.1, noticeable oscillations are observed, requiring a longer time to stabilize at the minimum value. This behavior improves when the learning rate is reduced to 0.01. Additionally, as expected, when the learning rate decreases, the number of k-iterations increases, indicating a slower convergence process. To highlight the high accuracy of the nonlocal continuous model compared to the discrete version, we have overlaid the discrete Adam simulations on the plot. This allows for a clear visualization of the remarkable similarity between both models.
@@
Figure 13 highlights the evolution of the first moment $m(t)$. For a learning rate of $\alpha=0.1$, $m(t)$ demonstrates a more pronounced oscillatory behavior before stabilizing around zero. When compared to the standard Adam optimizer’s convergence of the first moment $m_{k}$, both the nonlocal and standard Adam models show identical initial oscillatory patterns. With a reduced learning rate ($\alpha=0.01$), both models achieve smoother and more gradual convergence. The key difference is that, regardless of the learning rate, the curve for $\beta=0.0$ begins more smoothly in the nonlocal model, while the standard model shows a more abrupt start.
Figure 13: Convergence trajectories of $m(t)$ for the first-order nonlocal continuous Adam optimizer. The plot shows the evolution of the first moment $m(t)$ over time, scaled by the learning rate. With a higher learning rate ($\alpha=0.1$), the values of $m(t)$ exhibit certain oscillations, whereas a lower learning rate ($\alpha=0.01$) results in a smoother and slower convergence. To demonstrate the high level of accuracy of the nonlocal continuous model in replicating the discrete dynamics, the plot includes the discrete Adam simulations, allowing for a clear comparison between both approaches.
@@
Figure 14 shows the evolution of the second moment $v(t)$. For a higher learning rate ($\alpha=0.1$), the nonlocal continuous Adam model displays varying peak values depending on the parameter settings. Configurations such as $\beta_{1}=0.0$ and $\beta_{2}=0.999$ produce the highest peaks and the slowest decay. When compared to the standard Adam optimizer’s second moment $v_{k}$, both models follow similar trends, with distinct peak behaviors based on the parameter settings. Notably, the peaks for $\beta_{2}=0.999$ are generally higher in the nonlocal model, whereas for $\beta_{2}=0.99$, the peaks are slightly lower.
Figure 14: Convergence trajectories of $v(t)$ for the first-order nonlocal continuous Adam optimizer. The plot illustrates the convergence of the squared gradient moving average $v(t)$. At a higher learning rate ($\alpha=0.1$), the peaks of $v(t)$ vary according to the parameter settings, with $\beta_{1}=0.0$ and $\beta_{2}=0.999$ producing the most significant peak and the slowest rate of decay. When the learning rate is reduced to $0.01$, all curves display a more pronounced initial peak followed by a gradual decline. To enhance the visual comparison and highlight the accuracy of the nonlocal continuous model, the plot also includes the results of the discrete Adam optimizer, allowing the reader to directly observe the close alignment between the two methods.
@@
For a smaller learning rate ($\alpha=0.01$), the model shows an increase in the size of the peaks across all parameter settings, with a more pronounced bump for $\beta_{1}=0.0$, identical to the behavior observed in the standard case.

### v2

@@ 5 Numerical Simulations
In this section, we present numerical simulations [^8] All simulations can be reproduced from the repository available at GitHub. that validate the approximations proposed in the preceding sections. As shown, the simulations closely mirror the behavior of the discrete algorithms, confirming that the first-order integro-differential equations serve as accurate and reliable models for analyzing the dynamics of these optimization methods, both for convex and for nonconvex functions.
@@ 5.1 Numerical Method
To solve these equations, we employed the IDESolver method as described in Gelmi & Jorquera (2014), which we implemented from scratch, tailoring it to our specific requirements rather than relying on the existing Python module. See the pseudocode in Algorithm 4. Moreover, this implementation was carried out within the JAX framework in order to fully exploit GPU-based numerical computation. The GPU used was an NVIDIA T4.
@@
For the integral calculations, we employed a Gaussian quadrature method adapted for JAX with $n=1000$, instead of the quad method used in the original paper. This leads to a significant increase in calculation speed.
Additionally, we used the Euler method to solve the differential equation, rather than the RK45 method commonly employed in the original approach, to ensure consistency with Assumption 2. The (future) time increment term $\alpha$ of Propositions 1 and 2 is obtained using the interp1d interpolation function from the Interpax library.
@@
8:  Calculate the current global error $\text{new\_error}\leftarrow\|y_{\text{new}}-y_{\text{guess}}\|$
9:  if $\text{new\_error}>\text{error}$ then
@@
21:  Update $\text{error}\leftarrow\text{new\_error}$
@@ 5.2 Numerical Results
In this section, we present the numerical results obtained using the numerical method described previously and compare them with the outcomes derived from the original algorithms. All the figures are provided in Appendix B.
@@
5.2.1 Convex Function Simulations
We now begin with the simulations for the convex function $f(\theta)=(\theta-4)^{2}$, which has a unique and well-defined minimum at $\theta^{*}=4\,.$ We begin with the nonlocal model of AdaGrad, continue with RMSProp, and conclude with Adam.
@@ 5.2.1 Convex Function Simulations
Figure 1 shows the convergence of $\theta$ values using the nonlocal continuous AdaGrad method for minimizing the function $(\theta-4)^{2}$, evaluated at two learning rates: $0.1$ and $0.01$. This figure demonstrates that the nonlocal continuous AdaGrad method effectively replicates the optimization dynamics of the conventional AdaGrad method.
@@
In these simulations, AdaGrad is represented by $\xi=0$ in Theorems 1 and 2. Theorem 1 predicts the near-exponential approach of $\theta(t)$ to the minimizer observed in Fig. 1. In turn, Theorem 2 explains the behavior of the memory: with $\xi=0$ we only have $\dot{E}_{\rho}\leq-\,m\mu_{0}\,\Phi$, so $E_{\rho}$ is nonincreasing but not forced to decay exponentially, and the accumulated squared–gradient $G(t)$ is monotone increasing and converges to a finite $G_{\infty}$, as shown in Fig. 2. Consequently, the effective step $\mu(t)=1/\sqrt{G(t+\alpha)+\varepsilon}$ decreases and stabilizes at $1/\sqrt{G_{\infty}+\varepsilon}$, which explains the gradual flattening of the $\theta$–trajectories.
@@
For $\theta$, with a learning rate of 0.1, $\beta=0.99$ leads to faster convergence, while $\beta=0.0$ and $\beta=0.9$ result in slower convergence, with the latter exhibiting more linear behavior but showing noticeable oscillations. When the learning rate is reduced to 0.01, the overall convergence slows, and the oscillatory behavior for $\beta=0.0$ becomes less pronounced [^9] A slight oscillation is observed for $\beta=0.9$ and a learning rate of 0.01, induced by the numerical method, as the convergence error does not decrease below $1\times 10^{-4}$..
@@
When compared to the standard RMSProp optimizer, the nonlocal continuous variant exhibits a broadly similar convergence pattern, albeit with subtle differences. For instance, for $\beta=0.0$ and $\beta=0.9$, the nonlocal continuous model displays a slightly different dynamic than the discrete case, attaining marginally higher values. However, as the learning rate decreases, these differences gradually vanish, and the two models become nearly indistinguishable.
Finally, these simulations are represented by $\xi=\lambda$ in Theorems 1 and 2. RMSProp has a forgetting term $\lambda\,v$. The functional $E_{\rho}(t)=\Phi(t)+\rho\,v(t)$ therefore —Theorem 2— satisfies $\dot{E}_{\rho}\leq-\kappa_{0}E_{\rho}$, which implies exponential decay of $E_{\rho}$, $\Phi$ and $v$. This is exactly what Fig. 4 shows. In the memory plots (Fig. 4), $v(t)$ decays exponentially to 0, and the decay is slower as $\beta$ increases (since $\lambda=(1-\beta)/\alpha$ is smaller). In the $\theta$-trajectories (Fig. 3), Theorem 1 predicts the approach of $\theta(t)$ to the minimizer, as observed in Fig. 1.
@@
Figures 5, 6, and 7 illustrate the behavior of the first-order nonlocal continuous Adam optimization model by showing the convergence trajectories of $\theta(t)$, $m(t)$, and $v(t)$ over time for different parameter settings ($\beta_{1}$ and $\beta_{2}$) and learning rates ($\alpha=0.1$ and $\alpha=0.01$).
@@
Figure 6 highlights the evolution of the first moment $m(t)$. For a learning rate of $\alpha=0.1$, $m(t)$ demonstrates a more pronounced oscillatory behavior before stabilizing around zero. When compared to the standard Adam optimizer’s convergence of the first moment $m_{k}$, both the nonlocal and standard Adam models show identical initial oscillatory patterns. With a reduced learning rate ($\alpha=0.01$), both models achieve smoother and more gradual convergence.
@@
Figure 7 shows the evolution of the second moment $v(t)$. For a higher learning rate ($\alpha=0.1$), the nonlocal continuous Adam model displays varying peak values depending on the parameter settings. Configurations such as $\beta_{1}=0.0$ and $\beta_{2}=0.99$ produce the highest peaks and the slowest decay. When compared to the standard Adam optimizer’s second moment $v_{k}$, both models follow similar trends, with distinct peak behaviors based on the parameter settings. Notably, the peaks for $\beta_{1}=0.0,\,\,\beta_{2}=0.99\,$ and $\beta_{1}=0.9,\,\,\beta_{2}=0.99\,$ are generally higher in the nonlocal model.
@@
For a smaller learning rate ($\alpha=0.01$), the model shows an increase in the size of the peaks across all parameter settings, with a more pronounced bump for $\beta_{1}=0.0$ and $\beta_{2}=0.99$ or $\beta_{1}=0.9$ and $\beta_{2}=0.99$ , identical to the behavior observed in the standard case.
To conclude this section, let us analyze these plots in light of Theorem 3. Under the growth bound $C_{\alpha}<2\lambda_{1}$ and $\alpha^{*}>0$, we get $\|m(t)\|\!\to\!0$ as $t\to\infty$. Consequently, $\|\nabla f(\theta(t))\|\!\to\!0$, $v(t)\to 0$, and $\theta(t)\to\theta^{*}$. As shown, the figures confirm these theoretical predictions.
5.2.2 Nonconvex Function Simulations
Finally, we present the simulations for the non-convex function $f(\theta)=\frac{1}{4}(\theta^{2}-1)^{2}$, which has different local minima at $\theta=\pm 1$ and an unstable critical point (a local maximum) at 0. In this subsection, we restrict our discussion to the case $\theta=\pm 1$, since it is of particular interest how a change in the initial conditions may lead the system to one minimum or the other. For illustrations of the case $\theta=0$, please refer to Appendix A. All the figures are placed at Appendix B.
Before proceeding, we would like to make a brief observation regarding the position of the minima $\pm 1$ of the nonconvex function under study. Note that they are located as mirror images of each other, so it is to be expected that the simulations will exhibit mirror symmetry when comparing the negative and positive cases, as we shall see for AdaGrad, RMSProp, and Adam.
Nonlocal AdaGrad:
Figures 8(a), 8(b), 9(a), and 9(b) illustrate the behavior of the nonlocal continuous AdaGrad model compared with its discrete counterpart in a nonconvex setting, for two initial conditions of $\theta$ ($\theta_{0}=-0.1$ and $\theta_{0}=0.1$) and two learning rates ($\alpha=0.1$ and $\alpha=0.01$).
The trajectories in Figures 8(a) and 8(b) converge monotonically toward the stable minima ($\theta^{*}\approx-1$ and $\theta^{*}\approx 1$). For $\alpha=0.1$, the nonlocal model approaches the minimum a bit more slowly than the discrete version, particularly in the early stages, although both coincide in the long run. When the learning rate is reduced to $\alpha=0.01$, the two trajectories almost overlap, confirming the accuracy of the continuous model for smaller values of $\alpha$ and highlighting the system’s symmetry: starting at $\pm 0.1$ produces mirror profiles toward $\pm 1$, as expected.
The gradient plots in Figures 9(a) and 9(b) show a similar picture. With $\alpha=0.1$, the nonlocal model reaches a slightly higher plateau but follows a smoother transition, while for $\alpha=0.01$ the two curves are nearly indistinguishable across the whole time scale.
Let us discuss this figures through Lemma 11 and Theorems 4 and 5. In the figures with $\theta_{0}<0$ (Fig. 8(b)) and $\theta_{0}>0$ (Fig. 8(a)), the trajectories evolve within their respective basins and settle at limiting values. This is consistent with Lemma 11 ($\|\nabla f(\theta(t))\|\to 0$) and Theorem 4 (convergence to a critical point in $\mathcal{S}_{0}$). Furthermore, the near-exponential shape observed close to the limit is consistent with the PL case or with a KL exponent $\sigma=1/2$. For $\xi=0$ (Theorem 5, AdaGrad), the functional $E_{\rho}(t)=\Phi(t)+\rho\,G(t)$ is nonincreasing, but the dissipation inequality does not close as $\dot{E}_{\rho}\leq-\kappa E_{\rho}$; in particular, $G(t)$ is nondecreasing, as also occurs in the convex case. Therefore, no further decay rate for $G(t)$ is implied in this case, as illustrated in Fig. 9(a) or 9(b).
Nonlocal RMSProp:
Figures 10(a), 10(b), 11(a), and 11(b) compare the first–order nonlocal continuous RMSProp model with its discrete counterpart in the nonconvex case, for two initializations $\theta_{0}\in\{-0.1,\,0.1\}$.
Figures 10(a) and 10(b) show monotone convergence to the stable minima $\theta^{*}\approx-1$ and $\theta^{*}\approx 1$, respectively. With the larger learning rate $\alpha=0.1$, high momentum ($\beta=0.99$) produces pronounced overshoots and oscillations around the target, whereas $\beta=0$ and $\beta=0.9$ yield smoother trajectories that stabilize more quickly. Reducing the learning rate to $\alpha=0.01$ eliminates oscillations altogether: all settings converge smoothly, with $\beta=0.99$ converging the fastest, followed by $\beta=0.9$, and finally $\beta=0$. The dynamics are symmetric with respect to the sign of the initial condition, as we discussed at the beginning of this section.
On the other hand, figures 11(a) and 11(a) report the trajectories of $v(t)$. For $\alpha=0.1$, the nonlocal model produces smoother curves than the discrete scheme, particularly for $\beta=0$ and $\beta=0.99$. When $\beta=0$ or $0.9$, $v(t)$ exhibits a transient rise followed by decay, whereas for $\beta=0.99$ the values remain small and relatively flat—a behavior that does not occur in the discrete case, where at $t/\alpha\approx 5$ the curve rises noticeably. With $\alpha=0.01$, discrete and nonlocal curves nearly coincide for all $\beta$: for $\beta\in\{0,0.9\}$, they exhibit the characteristic bell-shaped peak (largest for $\beta=0$) before decaying to zero, while $\beta=0.99$ keeps $v(t)$ uniformly low with a slow decrease. As in the case of $\theta(t)$, the behavior is mirrored between $\theta_{0}=-0.1$ and $\theta_{0}=0.1$.
Once again, let us read the plots via Theorems 4, 5, and Lemma 11. We restrict the discussion to statements that follow directly from these results for RMSProp ($\xi=\lambda$). In the $v$-figures (e.g., Figs. 11(a)11(b)), $v(t)$ decays towards 0, in line with Theorem 5; the relative speed across $\beta$ agrees with the dependence of $\kappa$ on $\lambda$. In the $\theta$-panels (Figs. 10(a)10(b)), trajectories converge to a limiting value within $\mathcal{S}_{0}$, consistent with $\|\nabla f(\theta(t))\|\to 0$ (Lemma 11) and with the PL/KL conclusions of Theorem 4.
Nonlocal Adam:
Figures 12(a), 12(b), 13(a), and 13(b), as well as 14(a) and 14(b), compare the first–order nonlocal continuous Adam model with its discrete counterpart for both signs of the initialization $\theta_{0}=\pm 0.1$ in the nonconvex setting.
Figures 12(a) and 12(b) show a rapid approach to the stable minimizers ($\theta^{*}\approx\pm 1$). With the larger learning rate $\alpha=0.1$, settings with $\beta_{1}=0.9$ display overshoot and lightly damped oscillations around the limit, whereas $\beta_{1}=0$ stabilizes more quickly. The nonlocal trajectories follow the same trends (reduced oscillation amplitude) as the discrete ones. Lowering the learning rate to $\alpha=0.01$ removes instability: all parameter choices converge smoothly and almost perfectly overlap between the nonlocal and discrete models. The dynamics are symmetric with respect to the sign of $\theta_{0}$, as discussed earlier.
In Figures 13(a) and 13(b), $m(t)$ exhibits damped oscillations around zero for $\alpha=0.1$, with the most pronounced swings occurring for $\beta_{1}=0.9$ (especially with $\beta_{2}=0.99$ and $0.999$). For $\alpha=0.01$, all curves show a single valley followed by a gradual decay to zero; across the whole range the nonlocal and discrete trajectories are nearly indistinguishable. As expected, the profiles for positive and negative $\theta_{0}$ are mirror images.
Figures 14(a) and 14(b) show the evolution of the squared–gradient moving average $v(t)$. With $\alpha=0.1$, the peak height and decay rate depend strongly on the parameters: $\beta_{1}=0.9,\beta_{2}=0.99$ yields the largest transient peaks, while $\beta_{2}=0.999$ (for either $\beta_{1}$) produces much smaller, flatter curves due to heavier averaging. For $\alpha=0.01$, both models present the characteristic bell–shaped peak (largest for $\beta_{2}=0.99$) followed by a slow decay, and the nonlocal curves practically coincide with the discrete ones.
To conclude, let us analyze these figures via Theorem 6. What the figures show (and what the theorem certifies) is the following: (i) In the first–moment plots (Figs. 13(a), 13(b)), one observes transient behavior that damps to $\|m(t)\|\to 0$, exactly as guaranteed by Theorem 6. (ii) In the second–moment plots (Figs. 14(a), 14(b)), there is an initial peak followed by a decaying tail leading to $v(t)\to 0$, again in agreement with the theorem. (iii) In the $\theta$–trajectories (Figs. 12(a), 12(b)), the solutions stabilize at limiting values inside $\mathcal{S}_{0}$; this is consistent with $\|g(t)\|\to 0$.